OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 2, – 16), and x intercepts at x = -6 and x = 2. (Do not include the negative sign in your answer.) y = x2 + 4x − ___

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

If p and q are the roots, then the quadratic equation will be: y = a(x - p)(x - q) Here p = -6 and q = 2. Substitute in the above equation. You can find a knowing that the vertex lies on the curve. So put x = -2, y = -16 and solve for a.

OpenStudy (anonymous):

what about q

OpenStudy (anonymous):

Oops i see

OpenStudy (anonymous):

@ranga a equals 1

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

Thats right!?

OpenStudy (anonymous):

y = x2 + 4x −1

OpenStudy (ranga):

yes a = 1. The equations is: y = x^2 + 4x - 12

OpenStudy (anonymous):

the one goes there like that

OpenStudy (anonymous):

The equations is: y = x^2 + 4x - 12 so the answer is 12?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

I thought u said the answer was 1

OpenStudy (ranga):

No, the value of "a" = 1. With a = 1, y = a(x - p)(x - q) becomes y = (x + 6)(x - 2) = x^2 + 4x - 12 And the blank in the answer is 12.

OpenStudy (anonymous):

OHHH Thanks so much btw is that you in the picture

OpenStudy (ranga):

yes and two dogs.

OpenStudy (anonymous):

I love the picture I didn't even notice the dogs the lol

OpenStudy (ranga):

thanks.